OpenStudy (bringmethehorizon_):

WILL GIVE MEDAL!!!! Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 A: -96 B: 0 C: 24 D: 48

OpenStudy (cardboardbox):

The notation might be getting you confused. If I want g(10) that means, look at the definition of g(x) and replace all the x's with 10. So we have g(10) = 10 - 8 = 2

OpenStudy (bringmethehorizon_):

Solve that?

OpenStudy (cardboardbox):

Same thing with f. We can go nuts with this and do f( g(10) ). Don't freak out! This just means: First find out what g(10) is. We found it to be 2. Now we can just do f(2).

OpenStudy (bringmethehorizon_):

f(2)=10-8=2 ?

OpenStudy (cardboardbox):

Exactly right

OpenStudy (cardboardbox):

Wait no

OpenStudy (cardboardbox):

I think you might be having trouble understanding functions. Do you know what it means to square a number?

OpenStudy (bringmethehorizon_):

Yes I do

OpenStudy (cardboardbox):

Ok cool. Let's do an example with that.

OpenStudy (bringmethehorizon_):

Okayyy

OpenStudy (cardboardbox):

To square a number, you just times it by itself. I can describe this mathematically as a function: square(x) = x * x This says I have a function called "square", it takes in a number - and for the purposes of referring to this number, we call it x. In order to calculate the output of our square function, we multiply x by itself. So now how do we use it? Let's say I want to calculate 5 squared. I can use my square function! Let's feed it 5. So we write square(5) Now to calculate it, I look at what's inside the brackets, and it is 5. So x is now 5. To calculate the output, it's x*x, or with my input, it's just 5*5. We end up with square(5) = 25

OpenStudy (cardboardbox):

Your f and g functions are a bit more abstract than this. It's not quite apparent what they might be used for. But that's not a problem. We can treat them exactly like we did the square function.

OpenStudy (bringmethehorizon_):

Wait go back to the actual problem. the first would be 72 and the second is 2 @cardboardbox

OpenStudy (cardboardbox):

If by first you mean f(10) and second you mean g(10) then yes

OpenStudy (bringmethehorizon_):

Yes that's what I meant

OpenStudy (bringmethehorizon_):

What do I do next?

OpenStudy (cardboardbox):

Say we have another function "cube" which multiplies a number by itself and then again once more. You may (for reasons unknown) want to find the square of 3 cubed. We can write this as square( cube(3) ). So first we figure out what cube(3) is, and finally we can square that number to get our result.

OpenStudy (bringmethehorizon_):

Confused again

OpenStudy (bringmethehorizon_):

What do I have to cube

OpenStudy (cardboardbox):

Cube(3) means 3*3*3 which is 27

OpenStudy (cardboardbox):

In my example I am cubing 3, then squaring the result

OpenStudy (cardboardbox):

So square( cube(3) ) means, first figure out cube(3), which turns out to be 27. Then square 27 to get 729.

OpenStudy (bringmethehorizon_):

Is c the answer?

OpenStudy (cardboardbox):

How did you arrive at that

OpenStudy (bringmethehorizon_):

Or do you subtract everything from 729?

OpenStudy (cardboardbox):

729 was just my example

OpenStudy (bringmethehorizon_):

Oh, but I need to do the original problem @cardboardbox